Write down the molecular orbital configurations of N2 and NO. Calculate their bond orders. Which of these two is expected to have paramagnetic behavior?
1
Expert's answer
2020-02-17T07:06:33-0500

"N_2=

(\sigma_{1s})^2(\sigma^*_{1s})^2

(
σ
_{2s}
)^
2
(
σ^
*
_{2s}
)^
2
(
π
_{2px}
)^
2
(
π
_{2py}
)^2

(
σ
_{2pz}
)^
2
(
π
*_
{2px}
)
^0
(
π^
*_{2py})^0(\pi^*_{2pz})^0"


"bond \ order=\frac{1}{2}\times(10-4)=3"


"NO=N_2=( \sigma_{1
s}
)^2(\sigma^*_{1s})^2
(
σ
_{2s}
)^
2
(
σ^
*
_{2s}
)^
2
(
π
_{2px}
)^
2
(
π
_{2py}
)^2

(
σ
_{2pz}
)^
1
(
π
*_
{2px}
)
^0
(
π^
*_{2py})^0(\pi^*_{2pz})^0"


bond order ="\frac{1}{2}\times(9-4)=2.5"

"N_2" has not any unpaired electron so it is diamagnetic whereas NO has 1 unpaired electron so it is paramagnetic.